The Maroochy Chamber of Commerce is interested in determining the relationship between the number of fine days each year and the number of interstate and overseas tourists visiting the Sunshine Coast each year, measured in thousands. Using annual data from 1993 to 2005 (inclusive), the following model was estimated: Y = -3.96 + 0.56 X


                 Error sum of squares: 382.6        Sum of squares of X: 913.1


Determine the lower limit for the 90% confidence interval for the slope correct to two decimal places.
Answer #1

The number of data points, n = 2005-1993+1 = 13

Degrees of freedom, df = n-2 = 13-2 = 11

For 90% confidence interval, and df = 11, critical t-value = 1.796  

(critical t-value is taken from the table of critical values for given df and confidence level)

Slope = 0.56

Error sum of squares, SSE = 382.6

Standard error, se = sqrt(SSE/(n-2))

= sqrt(382.6/(13-2))

= 5.8976

90% confidence interval

Lower limit = slope - t * se

= 0.56 - 1.796 * 5.8976

= -10.032

Upper limit = slope + t * se

= 0.56 + 1.796 * 5.8976

= 11.152